# CNIMA: A Universal Evaluation Framework and Automated Approach for Assessing Second Language Dialogues

### v1

@@ 6 Experiments
We first replicate the ESL experiments of Gao et al. (2024) for CSL and compare their findings in Section 6.1. We then assess the performance of our proposed automated approach for SL assessment in Section 6.2.
@@ 6.1 Transferability of the Evaluation Framework from ESL to CSL
Based on the human-annotated micro-level features, Gao et al. (2024) convert them into normalised counts and use them as input features to train an LR, NB and RF classifier to predict macro-level interactivity labels (i.e. topic management, tone appropriateness, conversation opening and closing). They found: (1) strong prediction performance; (2) high-impact micro-level features that are common across all interactivity labels (by interpreting feature importance given by the trained classifiers); and (3) high-impact micro-level features that are specifically predictive for an interactivity label. We follow Gao et al. (2024) to compute the common and specific features for (2) and (3); details in Appendix A.10.
We replicate their experiments here, using our annotated CSL data (CNIMA), to see whether their findings transfer across languages.
@@ 6.2 Evaluation of Automated Pipeline
We now evaluate our automated 3-step approach to predicting the overall dialogue quality score and present the results in Table 7. In terms of model names, each component denotes the model used in a particular step, e.g. “BERT+LR+LR” means we use BERT for step 1 and LR for steps 2 and 3. Also, “GPT4” refers to GPT-4o. For brevity, we only include LR results for the classical models as they all have similar performances. In addition to our 3-step approach, we also include 2 baselines that predict the overall dialogue quality score directly based on the input dialogue: (1) fine-tuned BERT (“BERT (One-step)”); and (2) one-shot GPT-4o with instructions (“GPT4 (One-step)”). We also include a variation where the first step uses human-annotated micro-level features (e.g. “Human+LR+LR”) to understand how much performance degrades when substituting them with predicted features.
@@
When we look at the fully automated pipeline (bottom 4 rows in Table 7), “‘BERT+BERT+BERT” and “‘GPT4+GPT4+GPT4” perform very strongly (0.807 and 0.791), demonstrating that we have a fully automated system that can reliably assess the quality score of a dialogue. We also notice an interesting trend: When we use either BERT or GPT4 for predicting the micro-level features, LR (i.e. “BERT+LR+LR” and “GPT4+LR+LR”) performs very poorly for predicting the overall score, even though the span prediction performance of BERT and GPT4 for the micro-level features is no+t poor (we will revisit this in Section 6.3). This suggests that the classical machine learning models are less tolerant of noise. [^13] Note that LR does not use the dialogue as input, so any noise in the intermediate features (micro-level or macro-level) will have a bigger impact.
Taking all these results together, we show that when it comes to assessing SL dialogue quality, it is important to take a pipeline approach to predict important intermediate features (i.e. micro- and macro-level features) before predicting the overall quality score. This design also has another advantage: it is more interpretable because, given an overall score, we can also look at the intermediate results to understand what or where went wrong, providing more nuanced feedback that can benefit both teachers and second language students. Lastly, the strong performance of GPT4 (which is only marginally behind fine-tuned BERT) also has another important implication: we can adapt our SL assessment system to another language without requiring large-scale manually-annotated data by prompting LLMs.
@@ 6.3 Additional Analyses
To understand how macro-level interactivity labels contribute to the overall quality score, we present their feature importance as learned by LR, RF and NB in the “Human+LR+LR”, “Human+RF+RF” and “Human+NB+NB” models in Table 8. Topic management appears to be the most important predictor, followed by tone appropriateness. Conversation opening and closing, on the other hand, has relatively small weights. These observations are supported by Roever and Ikeda (2023), who similarly found that topic flow and tone are dominant factors that decide dialogue quality.
Next, we assess how well BERT and GPT4 perform for the first (micro-level feature span prediction) and the second step predictions (macro-level label classification); full results in AppendixA.5 Table14 and Table15 for BERT and GPT4 respectively. To summarise, for most of the 17 micro-level features (step 1), both BERT and GPT4 perform very well (average F1 for BERT and GPT4 is 0.89 and 0.81, respectively). That said, BERT and GPT4 struggle on two features related to non-factive verbs, and this can be attributed to the rarity of these features in the Chinese Roever and Ikeda (2023). Overall, these results show that the proposed micro-features by Gao et al. (2024) transfer very well from English to Chinese.
@@ 7 Conclusion
We develop CNIMA, a CSL dataset with 10K dialogues annotated based on an evaluation framework that assesses micro-level features, macro-level interactivity labels, and overall quality scores. As the evaluation framework is originally designed for ESL dialogues Gao et al. (2024), our first contribution tests how well the framework transfers to CSL dialogues. We found that the evaluation framework is robust across languages and further revealed universal and language-specific insights about the relationships between micro-level and macro-level features. Our second contribution is we propose an automated approach that predicts the micro-level features, macro-level labels and overall quality scores in sequence. We experimented with classical models, BERT and GPT-4o, and found that the BERT and GPT4-o-based systems perform very well in predicting the overall dialogue quality score. Our approach is interpretable, and the LLM-based variant can be easily adapted to another language for second language assessment without requiring annotated training data.
@@ 8 Limitations
The original speech data is transcribed manually into text. Technically we can use speech-to-text software to automate this, but since we’re working with second language conversations, this can result in much noise in the translated text. Ultimately, this does mean fully automating second language assessment using interactive conversations remains a challenging task.
The scope of the current paper is limited to CSL, but it would be equally interesting to see how the evaluation framework would work for other SL dialogues. Admittedly the models we experiment with are limited in terms of novelty, but we contend that is not the focus of our contribution in this paper.
@@
This study is conducted under the guidance of the ACL Code of Ethics. We manually filtered out potentially offensive content and removed all information related to the identification of annotators. The annotation protocol is approved under The University of Melbourne Human Ethics LNR as 2022-24988-32929-3.
@@ A.6 Prompts for GPT-4o
Example: SPK1: 那你是没有做吗?
SPK2: [的确], 我一点也没学 (SPK1: So you didn’t do it? SPK2: [Indeed], I haven’t learned at all.)
@@
Example: [SPK2 你既然有反驳的能力, 你还是自学吧]
(SPK2: [Since you have the self-learning skills, you shall teach yourself.])
@@ A.7 Prompts for GPT-4o Dialogue Overall Evaluation
Output Fields score: The score of the interactivity of the Chinese second language dialogue (1 to 5). rationale: The reason why and how the score is made.
Evaluation Criteria 5: Smooth and fluent daily communication, easy and pleasant. 4: Somewhat less fluent communication, but the communication purpose is achieved. 3: Slightly awkward communication, such as not being able to immediately understand the other person’s question with hesitation. 2: Overall communication is not fluent and awkward, but some parts can be mutually understood. 1: Unable to accurately achieve the communication purpose, awkward conversation, failed to talk throughout the conversation.

### v2

@@ 6 Experiments
We first replicate the ESL experiments of Gao et al. (2025) for CSL and compare their results in Section 6.1. We then assess the performance of our proposed automated approach for SL assessment in Section 6.2.
@@ 6.1 Transferability of the Evaluation Framework from ESL to CSL
For the human-annotated micro-level features, Gao et al. (2025) convert them into normalised counts and use them as input features to train an LR, NB and RF classifiers to predict macro-level interactivity labels. They found: (1) strong prediction performance; (2) high-impact micro-level features that are common across all interactivity labels (by interpreting feature importance given by the trained classifiers); and (3) high-impact micro-level features that are specifically predictive for an interactivity label. In the following, we replicate the experiments using our annotated CSL data (CNIMA), to see whether the results transfer across languages; details can be found in Appendix A.10.
@@ 6.2 Evaluation of Automated Pipeline
We now evaluate our automated 3-step approach to predicting the overall dialogue quality score as results in Table 7. Our pipeline predicts micro- and macro-level features and dialogue overall quality scores. In terms of model names, each component denotes the model used in a particular step, e.g. “BERT+LR+LR” means we use BERT for step 1 and LR for steps 2 and 3. Also, “GPT4” refers to GPT-4o. For brevity, we only include LR results for the classical models as they all have similar performances. In addition to our 3-step approach, we also include 2 baselines that predict the overall dialogue quality score directly based on the input dialogue: (1) fine-tuned BERT (“BERT (One-step)”); and (2) one-shot GPT-4o with instructions (“GPT4 (One-step)”). We also include a variation where the first step uses human-annotated micro-level features (e.g. “Human+LR+LR”) to understand how much performance degrades when substituting them with predicted features.
@@
When we look at the fully automated pipeline (bottom 4 rows in Table 7), “BERT+BERT+BERT” and “GPT4+GPT4+GPT4” perform very strongly (0.807 and 0.791), demonstrating that we have a fully automated system that can reliably assess the quality score of a dialogue. We also notice an interesting trend: When we use either BERT or GPT4 for predicting the micro-level features, LR (i.e. “BERT+LR+LR” and “GPT4+LR+LR”) performs very poorly for predicting the overall score, even though the span prediction performance of BERT and GPT4 for the micro-level features is not poor (we will revisit this in Section 6.3). This suggests that the classical machine learning models are less tolerant of noise. [^13] Note that LR does not use the dialogue as input, so any noise in the intermediate features (micro-level or macro-level) will have a bigger impact.
Taking all these results together, we show that when it comes to assessing SL dialogue quality, it is important to take a pipeline approach to predict important intermediate features (i.e. micro- and macro-level) before predicting the overall quality score. This design also has another advantage: it is more interpretable as given an overall score, we can also look at the intermediate results to understand what and where went wrong, providing more meaningful feedback that can benefit both teachers and second language students. Lastly, the strong performance of GPT4 also has another important implication: we can adapt our SL assessment system to another language without requiring large-scale manually-annotated data.
@@ 6.3 Additional Analyses
To understand how macro-level interactivity labels contribute to the overall quality score, we present their feature importance as learned by classical machine learning models in the “Human+LR+LR”, “Human+RF+RF” and “Human+NB+NB” models in Table 8. Topic management appears to be the most important predictor, followed by tone appropriateness. Conversation opening and closing, on the other hand, has relatively small weights. These observations are supported by Roever and Ikeda (2023), who similarly found that topic flow and tone are dominant in deciding interaction quality.
Next, we assess how well BERT and GPT4 perform for the first (micro-level feature span prediction) and the second-step predictions (macro-level label classification); full results in AppendixA.5 Table14 and 15 for BERT and GPT4 respectively. To summarise, for most of the 17 micro-level features (step 1), both BERT and GPT4 perform very well (average F1 for BERT and GPT4 is 0.89 and 0.81, respectively). That said, BERT and GPT4 struggle on two features related to non-factive verbs, and this can be attributed to the rarity of these features in the Chinese Roever and Ikeda (2023). Overall, these results show that the proposed micro-features transfer very well from English to Chinese.
@@ 7 Conclusion
We introduce CNIMA, a large-scale annotated dataset for Chinese-as-a-Second-Language (CSL) dialogues based on the evaluation framework introduced by Gao et al. (2025). This allows us to test the cross-lingual transferability of the framework, as it is originally designed for English. Our experiments demonstrate that the framework works well for CSL dialogues, suggesting that it is robust across languages and that it is a viable framework for evaluating any second-language conversations. Our analyses also reveal language-specific and language-universal relationships between micro-level linguistic features (such as backchannels) and macro-level interactivity features (such as topic management). Next, we propose an automated, interpretable approach that predicts the micro- and macro-level features and the overall quality of second-language dialogues. To this end, we explore a range of models, from classical machine learning models to LLMs. Our approach is interpretable because it does not only predict the overall dialogue quality score, but also reveals the key micro-level and macro-level features that contribute to the score. Our LLM-based approach performs very well, and as it does not require any labelled data for training, it can be adapted to other languages easily. Ultimately, our method paves the way for an interpretable and practical system for automatic second-language dialogue assessment.
@@ 8 Limitations
The original speech data is transcribed manually into text. Technically we can use speech-to-text software to automate this, but since we are working with second language conversations, this can result in much noise in the translated text. The scope of the current paper is limited to CSL, but it would be equally interesting to see how the evaluation framework would work for other SL dialogues, which calls on a joint contribution of the community effort on more SL language dialogue datasets, like Japanese second language datasets, Spanish second language datasets and etc.
@@
This study is conducted under the guidance of the ACL Code of Ethics. We manually filtered out potentially offensive content and removed all information related to the identification of annotators.
@@ A.6 Prompts for GPT-4o
Example: SPK1: 那你是没有做吗? SPK2: [的确], 我一点也没学 (SPK1: So you didn’t do it? SPK2: [Indeed], I haven’t learned at all.)
@@
Example: [我还是自学吧] (SPK2: [I shall teach myself.])
@@ A.7 Prompts for GPT-4o Dialogue Overall Evaluation
Output Fields score: The score of the interactivity of the Chinese second language dialogue (1 to 5).
rationale: The reason why and how the score is made.
Evaluation Criteria 5: Smooth and fluent daily communication, easy and pleasant.
4: Somewhat less fluent communication, but the communication purpose is achieved.
3: Slightly awkward communication, such as not being able to immediately understand the other person’s question with hesitation.
2: Overall communication is not fluent and awkward, but some parts can be mutually understood.
1: Unable to accurately achieve the communication purpose, awkward conversation, failed to talk throughout the conversation.